OpenStudy (anonymous):

find all trigonometric values of x with the given conditions 1. cosx= -15/17, sinx>0 2. tanx=-1, sinx<0 please help :)

OpenStudy (dumbcow):

set up a triangle, then you can find all 6 trig ratios rather easily if you remember SOH CAH TOA  now find missing side using pythorean thm: 17^2 - 15^2 = 8^2  hence sin x = 8/17 tan x = -8/15 the other 3 , sec,csc,cot are just the reciprocals sec x = -17/15 csc x = 17/8 cot = -15/8